OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, -3) in slope-intercept form

OpenStudy (anonymous):

Slope = \(\dfrac{rise}{run} = \dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

Then use the point-slope formula to write the equation. \((y-y_1)=m(x-x_1)\) Solve for y to get the slope-intercept form.